OpenStudy (anonymous):

The number of boys in ninth grade is 12 less than than twice the number of girls. There are 51 students in ninth grade. How many are girls?

OpenStudy (anonymous):

n=12-g n+g=51

OpenStudy (anonymous):

let n represent boys let g represent girls

OpenStudy (anonymous):

G = 21

OpenStudy (anonymous):

B + G = 51 , and B = 2G - 12 Now, replace the B in the first equation with its value from the second equation (2G - 12) + G = 51 Remove the brackets 2G - 12 + G = 51 Add 12 to both sides 2G + G = 51 + 12 3G = 63 Divide both sides by 3 G = 21 Now, to check if we're right, Replace G in the first equation B + 21 = 51 B = 30 If that's right then 30 = 2(21) - 12 30 = 42 - 12 Read more: http://wiki.answers.com/Q/The_number_of_boys_in_ninth_grade_is_12_less_than_twice_the_number_of_girls_There_are_51_students_in_ninth_grade_How_many_are_girls#ixzz1svgzE8Hv

OpenStudy (anonymous):

oh nvm n=12-2g

OpenStudy (anonymous):

So the answer to how many girls there are the answer is 21

OpenStudy (anonymous):

yes 21